OpenStudy (anonymous):

use properties of logarithms to expand the logarithmic expression as much as possible. log 11^(11x) is the answer x? I have no clue

OpenStudy (anonymous):

is this: \[\log_{11}11^{x}\] or\[\log11^{11x}\] ?

OpenStudy (anonymous):

The first one...

OpenStudy (anonymous):

ah, so you want to use the property: \[\log(a^{b}) = b*\log(a)\] so now we have: \[xlog_{11}11 = x(1) = x\]

OpenStudy (anonymous):

Or \[\log_{11} 11x\]

OpenStudy (anonymous):

Ooh, okay, I think I got it...